The partial pressure of Ch4(g) is 0.145 atm and that of O2(g) is 0.200 atm in a mixture of the two gases.

What is the mole fraction of each gas in the mixture?
1
Expert's answer
2017-11-27T11:36:06-0500
p_i/p_total =x_i=n_i/n_total , where x_i- mole fraction.
x_(〖CH〗_4 )= (0.145 atm)/((0.145+0.200) atm)×100%=42.0%; x_(O_2 )=(0.200 atm)/((0.145+0.200) atm)×100%=58.0%.

Answer: x_(〖CH〗_4 )=42.0%; x_(O_2 )=58.0%.